Present value of annuity​ payments)  The state​ lottery's million-dollar payout provides for ​$1.1 million to be paid in 20 installments of ​$55000 per payment. The first ​$55000 payment is made​ immediately, and the 19 remaining ​$55000 payments occur at the end of each of the next 19 years. If 12 percent is the discount​ rate, what is the present value of this stream of cash​ flows? If 24 percent is the discount​ rate, what is the present value of the cash​ flows? a.  If 12 percent is the discount​ rate, the present value of the annuity due is ​$ nothing. ​(Round to the nearest​ cent.)
Answer #1

PVAnnuity Due = c*((1-(1+ i)^(-n))/i)*(1 + i )
C = Cash flow per period
i = interest rate
n = number of payments
PV= 55000*((1-(1+ 12/100)^-20)/(12/100))*(1+12/100)
PV = 460117.73
PVAnnuity Due = c*((1-(1+ i)^(-n))/i)*(1 + i )
C = Cash flow per period
i = interest rate
n = number of payments
PV= 55000*((1-(1+ 24/100)^-20)/(24/100))*(1+24/100)
PV = 280319.51